OpenStudy (red3134):

Given: f(x) = x + 2 and h(x) = 1/ x - 1 2f(x) - h(x) = ?

OpenStudy (eliesaab):

2 f(x)= 2( x+2)=?

jhonyy9 (jhonyy9):

h(x) = (1/x) -1 or h(x) = 1/(x-1) ???

OpenStudy (red3134):

@jhonyy9 h(x) = 1/(x-1)

jhonyy9 (jhonyy9):

ok. - ty so can you calculi what wrote above eliesaab ?

OpenStudy (red3134):

@jhonyy9 Not really...

jhonyy9 (jhonyy9):

this will give you 2f(x) = ?

jhonyy9 (jhonyy9):

why ? f(x) = x+2 and this you need multiply it by 2 so will be 2(x+2) = ?

jhonyy9 (jhonyy9):

how you distribute this parentheses ?

OpenStudy (red3134):

@jhonyy9 would it be 2x + 4?

jhonyy9 (jhonyy9):

exactly - very nice

jhonyy9 (jhonyy9):

so and now 2f(x) -h(x) = ?

OpenStudy (red3134):

@jhonyy9 (2x + 4) - 1/ (x - 1) ?

jhonyy9 (jhonyy9):

yes and make the calculi

jhonyy9 (jhonyy9):

a-1/b = ?

jhonyy9 (jhonyy9):

how you subtract from one integer a fraction ? a-(1/b) = ?

jhonyy9 (jhonyy9):

how many will be the common denominator ?

jhonyy9 (jhonyy9):

in case of my example the common denominator will be b so this mean that you need multiply the a by b and will get (a*b -1)/b ok. ? can use this example to solve your exercise ?

jhonyy9 (jhonyy9):

what is b in case of your exercise ?

jhonyy9 (jhonyy9):

please use this example - so b in case of your exercise will be (x-1)

OpenStudy (red3134):

Would a be 2x + 4?

OpenStudy (red3134):

@jhonyy9 And 2x + 4 multiplied by x - 1 would be 2x - 4?

jhonyy9 (jhonyy9):

no 2x+4 - ( 1/(x-1) ) = ?

jhonyy9 (jhonyy9):

(2x+4)(x-1) = ?

jhonyy9 (jhonyy9):

= 2x*x +2x*(-1) +4*x +4*(-1) = ?

OpenStudy (red3134):

\[2x ^{2}+ 2x - 4\]

jhonyy9 (jhonyy9):

yes but dont forget the denominator 2f(x) -h(x) = ?

jhonyy9 (jhonyy9):

looke please in case of my example the common denominator will be b so this mean that you need multiply the a by b and will get (a*b -1)/b ok. ? can use this example to solve your exercise ?

jhonyy9 (jhonyy9):

this what you ve got above is just the part of a*b

jhonyy9 (jhonyy9):

do you understand it now ?

jhonyy9 (jhonyy9):

but you need getting (a*b -1)/b

jhonyy9 (jhonyy9):

so will be 2f(x) -h(x) = 2x+4 - (1/(x-1)) = [(2x+4)(x-1)-1]/(x-1) = ?

jhonyy9 (jhonyy9):

substitute there in place of (2x+4)(x-1) this above got value

jhonyy9 (jhonyy9):

= (2x^2 +2x-4 -1)/(x-1) = ?

jhonyy9 (jhonyy9):

-4-1 = ? this is the last step

OpenStudy (red3134):

Thank you for your help! I think I get it now :)

jhonyy9 (jhonyy9):

ok. good luck yw